OpenStudy (mony01):

What are the equilibrium points? dR/dt=R(1200-2R-3S) dS/dt=S(500-R-S)

OpenStudy (gorv):

when derivative equal=0 that will be equilibrium

OpenStudy (gorv):

R(1200-2R-3S)=0 >>>> 1200-2R-3S=0 S(500-R-S)=0 >>>>500-R-S=0

OpenStudy (mony01):

do i solve for R or S?

OpenStudy (mony01):

@gorv